Solve 5x + 2y = 7
10x + 4y = 14
Please show work

Answer by jim_thompson5910(35256)

Hint: Multiply everything in equation #1 by -2 to get -10x - 4y = -14

So we now have
-10x - 4y = -14
10x + 4y = 14

Add the equations. What do you get?